Graph the parabola
y=(x+1)^2-1
I found the vertex to be (-1,-1)
axis is -1 and a=1
But I don't remember how to find the points.

Found 2 solutions by Fombitz, stanbon:
Answer by Fombitz(32388)

Which points?
.
.
.
Intercepts?
Set x=0 solve for y.
Set y=0 solve for x.
.
.
.
Points on the curve?
Choose points equidistant from x=-1, solve for y.
Example, if x=0=-2, y=0

Answer by stanbon(75887)

Graph the parabola
y=(x+1)^2-1
I found the vertex to be (-1,-1)
axis is -1 and a=1
But I don't remember how to find the points.
---------------
You need three points to determine a parabola.
You have the vertex at (-1,-1)
The y-intercept ?
Let x = 0, then y = (0+1)^2-1 = 0
That gives you point (0,0)
------
The x-intercept ?
Let y = 0 and solve for "x" to get
(0,0) and (-2,0)
====
Cheers,
Stan H.